Calculate the potential of Ag-Zn cell at 298 K if the concentration of Ag+ and Zn2+ are 5.2 x 10-6 M and 1.3 x 10-3 M respectively. Eo of the cell at 298 K is 1.56 V. Write the cell representation and reactions. Calculate the change in free energy ∆G for the reduction of 1 mole of Ag+.
1
Expert's answer
2019-09-10T03:47:20-0400

Cell representation of this cell:

"Zn | Zn^{+2} || Ag^{+1} | Ag"

Cell reactions:

"Zn \to Zn^{+2}+2e^-"

"Ag^++e^- \to Ag"

Net reaction:"Zn +2Ag^+ \to Zn^{+2} +2Ag"

Number of electrons used"(n)=2"

"E_{cell}=E_0-\frac{0.0591}{n} \space log_{10} \frac{[Zn^{+2}]}{[Ag^+]^2}"

"E_{cell}=1.56-\frac{0.0591}{2} \space log_{10} \frac{1.3 \times 10^{-3}}{(5.2 \times 10^{-6})^2}"

"E_{cell}=1.56-0.227=1.33 V"

Gibbs energy change"=nFE=2 \times 96500 \times 1.33=256,690 J=256.7 KJ"